# Quantum Many-Body Scarring in $2+1$D Gauge Theories with Dynamical Matter

### v1

@@ Abstract
Introduction.—QMBS is an exciting paradigm of ergodicity breaking in isolated quantum many-body models that are expected to thermalize Bernien et al. (2017); Moudgalya et al. (2018a); Zhao et al. (2020); Jepsen et al. (2022); Serbyn et al. (2021); Moudgalya et al. (2022); Chandran et al. (2023). Despite being ergodic, certain models host special nonthermal scar eigenstates that are roughly equally spaced in energy over the whole spectrum Turner et al. (2018); Schecter and Iadecola (2019), and exhibit anomalously low bipartite entanglement entropy Moudgalya et al. (2018b); Lin and Motrunich (2019). Upon preparing an initial state with a high overlap with these scar eigenstates, the system avoids thermalization and its dynamics exhibits long-lived coherent oscillations lasting well beyond all relevant timescales, with the time-evolved wave function undergoing persistent periodic revivals Turner et al. (2018); Ho et al. (2019). QMBS is of great importance in investigations of the Eigenstate Thermalization Hypothesis (ETH) Deutsch (1991); Srednicki (1994); D’Alessio et al. (2016); Deutsch (2018), as it facilitates violations of the latter through novel mechanisms based on spectrum-generating algebras Mark et al. (2020); Moudgalya et al. (2020); O’Dea et al. (2020); Pakrouski et al. (2020) and nonthermal-eigenstate embedding Shiraishi and Mori (2017). QMBS has also been the subject of various experiments in both analog and digital quantum simulators Bernien et al. (2017); Bluvstein et al. (2021, 2022); Su et al. (2023); Zhang et al. (2023); Dong et al. (2023).
@@
QMBS also has an intimate connection to lattice gauges theories (LGTs) Surace et al. (2020); Iadecola and Schecter (2020); Aramthottil et al. (2022); Biswas et al. (2022); Desaules et al. (2023a, b); Halimeh et al. (2023a), which are interacting quantum many-body models hosting gauge symmetries that enforce an intrinsic relation between the local distribution of matter and the allowed corresponding configurations of the gauge fields Weinberg (1995); Peskin and Schroeder (2018). Indeed, the quantum Ising-like model realized in a Rydberg setup that produced the first instance of QMBS Bernien et al. (2017) can be effectively described by the PXP model, which Surace et al. Surace et al. (2020) have shown to map exactly onto the $1{+}1$D spin-$1/2$ $\mathrm{U}(1)$ QLM Chandrasekharan and Wiese (1997); Wiese (2013). The spin-$1/2$ $\mathrm{U}(1)$ QLM is a formulation of the Schwinger model in which the gauge and electric fields are represented by spin-$1/2$ operators. This truncation has facilitated experimental feasibility in the realization of large-scale quantum simulators of the spin-$1/2$ $\mathrm{U}(1)$ QLM Yang et al. (2020); Zhou et al. (2022), while still capturing salient features of the Schwinger model such as Coleman’s phase transition Coleman (1976). Given the ongoing strong drive of realizing quantum simulators of LGTs Dalmonte and Montangero (2016); Bañuls et al. (2020); Zohar et al. (2015); Alexeev et al. (2021); Aidelsburger et al. (2022); Zohar (2022); Klco et al. (2021); Bauer et al. (2023); Meglio et al. (2023); Halimeh et al. (2023b); Cheng and Zhai (2024), it is important to fully understand the connection between QMBS and LGTs both from a fundamental point of view and also to facilitate experimental investigations.
@@
In particular, it is interesting to see how QMBS behaves in $2{+}1$D LGTs. Indeed, it is known that nonergodic features in $1{+}1$D interacting models tend to vanish in higher spatial dimensions $d{>}1$, with many-body localization Basko et al. (2006); Gornyi et al. (2005); Nandkishore and Huse (2015); Abanin et al. (2019) being a prime example. How the weak ergodicity-breaking mechanism of QMBS fares for $d{=}2$ is a question that has recently received some attention in the context of the PXP model Michailidis et al. (2020); Lin et al. (2020) and helix states in the XXZ model Jepsen et al. (2022), in addition to other quasi-$2{+}1$D systems Zhang et al. (2023); Dong et al. (2023). Nevertheless, scarred models in $d{>}1$ are far and few in between compared to their counterparts in $d{=}1$ Serbyn et al. (2021); Moudgalya et al. (2022); Chandran et al. (2023). This further motivates adding to the collection of such models by investigating LGTs with dynamical matter in higher spatial dimensions.
@@
It is important to emphasize that our work does not rule out other QMBS regimes that may still persist in $2+1$D but that we have not found. Indeed, our goal in this work was to investigate the fate of QMBS regimes discovered in the $1+1$D $\mathrm{U}(1)$ QLM and see how well they fare in $2+1$D. It was also our intention to find out how the statistics of the matter degrees of freedom affects QMBS, which, as we show, happens in a profound way. One avenue for future work is to study the fate of the QMBS regime we find in $2+1$D for higher-level representations ($S>1/2$) of the electric and gauge fields. In $1+1$D, it has been shown that scarring persists and is robust for $S>1/2$ Desaules et al. (2023a, b). This is important to assess the fate of QMBS in the limit of $2+1$D lattice QED.
@@ Acknowledgements.
Note.—During the final stages of preparing our manuscript, we became aware of another work Bud on quantum many-body scars for arbitrary integer spin in $2+1$D pure Abelian gauge theories. This work will appear in the same arXiv listing as ours. Acknowledgments.—The authors are grateful to Jean-Yves Desaules and Zlatko Papić for fruitful discussions. This work is supported by the Emmy Noether Programme of the German Research Foundation (DFG) under grant no. HA 8206/1-1. I.P.M. acknowledges funding from the National Science and Technology Council (NSTC) Grant No. 122-2811-M-007-044. Numerical simulations were performed on The University of Queensland’s School of Mathematics and Physics Core Computing Facility getafix.

### v2

@@ Abstract
Introduction.—QMBS is an exciting paradigm of ergodicity breaking in isolated quantum many-body models that are expected to thermalize Bernien et al. (2017); Moudgalya et al. (2018a); Zhao et al. (2020); Jepsen et al. (2022); Serbyn et al. (2021); Moudgalya et al. (2022); Chandran et al. (2023). Despite being ergodic, certain models host special nonthermal scar eigenstates that are roughly equally spaced in energy over the whole spectrum Turner et al. (2018); Schecter and Iadecola (2019), and exhibit anomalously low bipartite entanglement entropy Moudgalya et al. (2018b); Lin and Motrunich (2019). Upon preparing an initial state with a high overlap with these scar eigenstates, the system avoids thermalization and its dynamics exhibits long-lived coherent oscillations lasting well beyond all relevant timescales, with the time-evolved wave function undergoing persistent periodic revivals Turner et al. (2018); Ho et al. (2019). QMBS is of great importance in investigations of the Eigenstate Thermalization Hypothesis (ETH) Deutsch (1991); Srednicki (1994); D’Alessio et al. (2016); Deutsch (2018), as it facilitates violations of the latter through novel mechanisms based on spectrum-generating algebras Mark et al. (2020); Moudgalya et al. (2020); O’Dea et al. (2020); Pakrouski et al. (2020) and nonthermal-eigenstate embedding Shiraishi and Mori (2017). QMBS has also been the subject of various experiments in both analog and digital quantum simulators Bernien et al. (2017); Bluvstein et al. (2021, 2022); Su et al. (2023); Zhang et al. (2023a); Dong et al. (2023).
@@
QMBS also has an intimate connection to lattice gauges theories (LGTs) Surace et al. (2020); Iadecola and Schecter (2020); Aramthottil et al. (2022); Biswas et al. (2022); Desaules et al. (2023a, b); Halimeh et al. (2023a), which are interacting quantum many-body models hosting gauge symmetries that enforce an intrinsic relation between the local distribution of matter and the allowed corresponding configurations of the gauge fields Weinberg (1995); Peskin and Schroeder (2018). Indeed, the quantum Ising-like model realized in a Rydberg setup that produced the first instance of QMBS Bernien et al. (2017) can be effectively described by the PXP model, which Surace et al. Surace et al. (2020) have shown to map exactly onto the $1{+}1$D spin-$1/2$ $\mathrm{U}(1)$ QLM Chandrasekharan and Wiese (1997); Wiese (2013). The spin-$1/2$ $\mathrm{U}(1)$ QLM is a formulation of the Schwinger model in which the gauge and electric fields are represented by spin-$1/2$ operators. This truncation has facilitated experimental feasibility in the realization of large-scale quantum simulators of the spin-$1/2$ $\mathrm{U}(1)$ QLM Yang et al. (2020); Zhou et al. (2022); Su et al. (2023); Zhang et al. (2023b), with proposed algorithms for digital platforms Barros et al. (2024), while still capturing salient features of the Schwinger model such as Coleman’s phase transition Coleman (1976). Given the ongoing strong drive of realizing quantum simulators of LGTs Dalmonte and Montangero (2016); Bañuls et al. (2020); Zohar et al. (2015); Alexeev et al. (2021); Aidelsburger et al. (2022); Zohar (2022); Klco et al. (2021); Bauer et al. (2023); Meglio et al. (2023); Halimeh et al. (2023b); Cheng and Zhai (2024), it is important to fully understand the connection between QMBS and LGTs both from a fundamental point of view and also to facilitate experimental investigations.
@@
In particular, it is interesting to see how QMBS behaves in $2{+}1$D LGTs. Indeed, it is known that nonergodic features in $1{+}1$D interacting models tend to vanish in higher spatial dimensions $d{>}1$, with many-body localization Basko et al. (2006); Gornyi et al. (2005); Nandkishore and Huse (2015); Abanin et al. (2019) being a prime example. How the weak ergodicity-breaking mechanism of QMBS fares for $d{=}2$ is a question that has recently received some attention in the context of the PXP model Michailidis et al. (2020); Lin et al. (2020) and helix states in the XXZ model Jepsen et al. (2022), in addition to other quasi-$2{+}1$D systems Zhang et al. (2023a); Dong et al. (2023). Nevertheless, scarred models in $d{>}1$ are far and few in between compared to their counterparts in $d{=}1$ Serbyn et al. (2021); Moudgalya et al. (2022); Chandran et al. (2023). This further motivates adding to the collection of such models by investigating LGTs with dynamical matter in higher spatial dimensions.
@@
It is important to emphasize that our work does not rule out other QMBS regimes that may still persist in $2+1$D but that we have not found. Indeed, our goal in this work was to investigate the fate of QMBS regimes discovered in the $1+1$D $\mathrm{U}(1)$ QLM and see how well they fare in $2+1$D. It was also our intention to find out how the statistics of the matter degrees of freedom affects QMBS, which, as we show, happens in a profound way. One avenue for future work is to study the fate of the QMBS regime we find in $2+1$D for higher-level representations ($S>1/2$) of the electric and gauge fields. In $1+1$D, it has been shown that scarring persists and is robust for $S>1/2$ Desaules et al. (2023a, b). This is important to assess the fate of QMBS in the limit of $2+1$D lattice QED, and a recent proposal may allow experimental observation in a quantum simulator of the spin-$1$ $\mathrm{U}(1)$ QLM Osborne et al. (2023). A second venue involves adding a topological $\theta$-term Halimeh et al. (2022); Cheng et al. (2022) and studying its interplay with scarring in $2+1$D, which has been shown to lead to a plethora of nonergodic behavior is $1+1$D Desaules et al. (2024).
@@ Acknowledgements.
Note.—During the final stages of preparing our manuscript, we became aware of another work Budde et al. (2024) on quantum many-body scars for arbitrary integer spin in $2+1$D pure Abelian gauge theories. This work will appear in the same arXiv listing as ours. Acknowledgments.—The authors are grateful to Jean-Yves Desaules and Zlatko Papić for fruitful discussions. This work is supported by the Emmy Noether Programme of the German Research Foundation (DFG) under grant no. HA 8206/1-1. I.P.M. acknowledges funding from the National Science and Technology Council (NSTC) Grant No. 122-2811-M-007-044. Numerical simulations were performed on The University of Queensland’s School of Mathematics and Physics Core Computing Facility getafix.